OpenStudy (geekfromthefutur):

plz help Medal and fan rewarded (: Maria is comparing the prices of two window cleaning companies. Company A charges $6 per window and an additional $12 as service charges. Company B charges $5 per window and an additional $15 as service charges.

OpenStudy (geekfromthefutur):

Part A: Write equations to represent Company A and Company B's total charges for cleaning a certain number of windows. Define the variable used in the equations. (4 points) Part B: Which company would charge less for cleaning 8 windows? Justify your answer. (3 points) Part C: How much money is saved by using the services of Company B instead of Company A to clean 6 windows? (3 points)

OpenStudy (anonymous):

Part A: y=6x+12 and y=5x+15 The variable used in both equations is x

OpenStudy (geekfromthefutur):

thx (:

OpenStudy (geekfromthefutur):

can u help me with the other 2 lolz

OpenStudy (anonymous):

For Part B you would just plug in 8 for x and the amount for company A would be y=6(8)+12 and the amount for company B would be 5(8)+15 and whichever one of equations has the highest answer is NOT the answer

OpenStudy (anonymous):

For part C plug in 6 for x in both equations and find the difference between the answers by subtracting the smallest number from the greatest number

OpenStudy (geekfromthefutur):

oh ok thank you :3

OpenStudy (geekfromthefutur):

@Drsuz98

OpenStudy (geekfromthefutur):

part b

OpenStudy (anonymous):

Okay the answer to y=6(8)+12 is 60 and the answer to y=5(8)+15 is 55

OpenStudy (anonymous):

so the answer to part B is company B

OpenStudy (geekfromthefutur):

ok thank you soooo much your a life saver (: can u helpme with part see i have alittle bit of knowing how to work it out but im getting stuck :/

OpenStudy (anonymous):

Yeah waht r u stuck om

OpenStudy (geekfromthefutur):

im stuck on which one would be the right answer but im thinken it company A

OpenStudy (anonymous):

okay what were ur two answers that u got fot the equations

OpenStudy (anonymous):

@geekfromthefutur

OpenStudy (geekfromthefutur):

sorry i was in the kitchen

OpenStudy (anonymous):

cool so what r they